Text Size:
Smaller Text Normal Text Larger Text

Facebook page

Twitter profile

YouTube channel

Instructions from the PI

Dear NOT observer and staff.

This is a set of instructions for NOT monitoring programme P68-509 (PI:HÃ¥kon Dahle), scheduled for service observations. The purpose is to monitor three of the five currently known gravitationally lensed quasars where the lens is a galaxy cluster. The main purpose of the observations is to use the light curves to make time delay measurements between multiple lensed quasar images in order to estimate the value of the Hubble constant. We will also determine the level of microlensing in each lens system. The target information and the observing strategy are reported below.

Please let us know whether you carried out these observations.

Thank you!

Best regards,
Håkon Dahle


For the "monitoring" observations use the script indicated by the OB.


For calibration we request:

- A set of three twilight flats for the SDSS g' filter
- alfosc-calibs

Available scripts

Go to the directory
  cd ~/scripts/68-509

For these observations use the appropriate script for the date, complie the OB if needed.


CJ0219+2156 13 minutes 16 epochs

CJ0335-1927 20 minutes 15 epochs

CJ0542-2125 20 minutes 17 epochs

J0909+44, 20 minutes 19 epochs

CJ1153+0756 35 minutes 15 epochs

J1326+48 13 minutes 15 epochs

J2222+2745 20 minutes 6 epochs

Date Time (UT) Duration Comments
29-03 (EAT-15m)-21:25 00:20 epoch18 SDSSJ0542-2125
31-03 (EAT-15m)-01:35 00:20 epoch20 SDSSJ0909+4449
31-03 (EAT-15m)-01:35 00:35 epoch15 SDSSJ1153+0755
31-03 21:00-01:35 00:13 epoch16 SDSSJ1326+4806
06-04 (EAT-15m)-02:15 00:20 SDSS J0909+4449 epoch1
06-04 (EAT-15m)-03:15 00:35 SDSS J1153+0755 epoch1
06-04 (EAT-15m)-(SAT+15m) 00:13 SDSS J1326+4806 epoch1
11-04 23:10-03:30 00:35 SDSS J1153+0755 epoch2
11-04 23:10-01:50 00:20 SDSS J0909+4449 epoch2
14-04 02:00-(SAT+15m) 00:13 SDSS J1326+4806 epoch2
19-04 04:55-(SAT+15m) 00:20 SDSS J2222+2745 epoch1
26-04 (EAT-15m)-22:30 00:20 SDSS J0909+4449 epoch3
28-04 (EAT-15m)-00:30 00:35 SDSS J1153+0755 epoch3
28-04 (EAT-15m)-00:30 00:13 SDSS J1326+4806 epoch3
02-05 (EAT-15m)-00:15 00:20 epoch4 SDSSJ0909+4449
02-05 (EAT-15m)-02:00 00:35 epoch4 SDSSJ1153+0755
04-05 (EAT-15m)-04:45 00:13 epoch4 SDSSJ1326+4806
09-05 (EAT-15m)-00:10 00:20 epoch5 SDSSJ0909+4449
09-05 (EAT-15m)-01:30 00:35 epoch5 SDSSJ1153+0755
13-05 01:20-04:15 00:13 epoch5 SDSSJ1326+4806
13-05 03:30-(SAT+15m) 00:20 epoch2 SDSSJ2222+2745
25-05 (EAT-15m)-22:25 00:20 epoch6 SDSSJ0909+4449
27-05 (EAT-15m)-00:10 00:13 epoch6 SDSSJ1326+4806
28-05 (EAT-15m)-00:10 00:35 epoch6 SDSSJ1153+0755
03-06 (EAT-15m)-22:30 00:20 epoch7 SDSSJ0909+4449
03-06 (EAT-15m)-02:45 00:13 epoch7 SDSSJ1326+4806
03-06 (EAT-15m)-23:45 00:35 epoch7 SDSSJ1153+0755
08-06 (EAT-15m)-23:30 00:35 epoch8 SDSSJ1153+0755
08-06 (EAT-15m)-22:15 00:20 epoch8 SDSSJ0909+4449
08-06 (EAT-15m)-02:30 00:13 epoch8 SDSSJ1326+4806
16-06 02:45-(SAT+15m) 00:20 epoch3 SDSSJ2222+2745
25-06 (EAT-15m)-23:30 00:13 epoch9 SDSSJ1326+4806
25-06 (EAT-15m)-22:45 00:35 epoch9 SDSSJ1153+0755

Program Tracking
Proposal 68-509
Semester P69: 20240401 - 20241001
Principal Investigator Hakon Dahle
Programme Cluster-lensed quasars as probes of H0
Instrument ALFOSC
Time Allocation 8x20m J0909, 15x13m J1326, 10x13m CJ0219, 6x20m CJ0335, 3x20m CJ0542, 9x35m CJ1153, 6x20m J2222
Total Nights 18.3h
Type Mon
Notes Fixed dates to be scheduled
Total Time Spent 06:53 (hh:mm)
Proposal Date Time Type Targets Files Comments Status
68-509 2024-04-06 00:45 Mon SDSSJ0909, SDSSJ1326, COOLJ1153 ALHd060175 - ALHd060177, ALHd060186, ALHd060206 - ALHd060208 SDSSJ0909 and SDSSJ1326 were observed with the given requirements. COOLJ1153 was also tried when the seeing looked good enough. After the first exposure, however, the seeing appeared to be ~1.5''. The script was stopped (the first exposure file can be checked) and no other time windows for the seeing requirement came later on to attempts the target again. Observed
68-509 2024-04-11 00:00 Mon Unable to observe any targets due to seeing conditions beyond defined constraints. Failed
68-509 2024-04-14 00:11 Mon SDSSJ1326 ALHd140090 - ALHd140092 The target SDSSJ1326 was observed when seeing was ~1.9'' after the Moon had gotten below horizon. Observed
68-509 2024-04-19 00:00 Mon Seeing worse than 1.3 and clashing with a hard ToO Failed
68-509 2024-04-26 00:18 Mon SDSSJ0909 ALHd260263 - ALHd260265 Observed in clear skies with seeing around 0.8" Observed
68-509 2024-04-28 00:00 Mon Failed because telescope could not be opened due to high humidity and subzero temperatures. Failed
68-509 2024-05-02 00:53 Mon SDSSJ0909, COOLJ1153 ALHe020152 - ALHe020154, ALHe020176 - ALHe020178 Observed in clear skies with seeing around 0.8" Observed
68-509 2024-05-04 00:12 Mon SDSSJ1326 ALHe040078 - ALHe040080 Observed in thin clouds with seeing around 0.9" Observed
68-509 2024-05-09 00:51 Mon SDSSJ0909, COOLJ1153 ALHe090160 - ALHe090162, ALHe090207 - ALHe090209 Both targets SDSSJ0909, COOLJ1153 observed through thin clouds and clear skies respectively. Observed
68-509 2024-05-13 00:31 Mon SDSSJ1326, SDSSJ2222 ALHe130109 - ALHe130111,ALHe130129 - ALHe130131 Targets observed with clear skies, 1" - 2" seeing. Observed
68-509 2024-05-25 00:00 Mon none Could not observe targets because of technical problems with ALFOSC. Failed
68-509 2024-05-27 00:11 Mon SDSSJ1326 ALHe270091 - ALHe270093 Observed in clear skies with seeing around 0.8" Observed
68-509 2024-05-28 00:34 Mon COOLJ1153 ALHe280162 - ALHe280164 Observed in clear skies with seeing around 1.0" Observed
68-509 2024-06-03 01:04 Mon SDSSJ0909, COOLJ1153, SDSSJ1326 ALHf030135 - ALHf030137, ALHf030142 - ALHf030144, ALHf030183 - ALHf030185 Clear skies and a seeing of less than 1 arcsec Observed
68-509 2024-06-08 01:02 Mon SDSSJ0909, COOLJ1153, SDSSJ1326 ALHf080164 - ALHf080166, ALHf080175 - ALHf080177, ALHf080185 - ALHf080187 All requested monitor targets observed. Clear skies, sub-arcsec seeing. Observed
68-509 2024-06-16 00:21 Mon SDSSJ2222 ALHf160104 - ALHf160106 Target observed with clear skies and subarcsec seeing Observed


The above information is collected from the ToO/Mon reports that have been created by the observer.

Should any of the observations have been carried out using the OB system, you will find additional information HERE

Back to top Last modified: May 24 2024